OpenStudy (anonymous):

What is the approximate circumference of a circle with a radius of 20 inches? Use π ≈ 3.14. A. 251.2 inches B. 62.8 inches C. 120 inches D. 125.6 inches

OpenStudy (anonymous):

@patricialarios

OpenStudy (anonymous):

It is D did you want the explination

OpenStudy (anonymous):

Nah i trust you :D

OpenStudy (anonymous):

It is because a radius is only half a diameter so you want to add 20 more to that 20 so make it a full diamerter the times it by 3.14 and it equals D lol

OpenStudy (anonymous):

Well thanks for the explanation~!

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

anymore?

OpenStudy (anonymous):

A circular lid has a diameter of 12 inches. What is the exact area of the lid? A. 12π square inches B. 24π square inches C. 36π square inches D. 6π square inches

OpenStudy (anonymous):

The question marks are the pi symbol

OpenStudy (anonymous):

Well Area=3.14*Radius^2 for a circle so Area= 3.14*6*6 Area=3.14*36 SO answer is C 36Pi

OpenStudy (anonymous):

Thannnnnk you~

OpenStudy (anonymous):

Welcome! if you have anymore can you make a new post LOL too many comments here

OpenStudy (anonymous):

Okay haha

OpenStudy (anonymous):

what math is this? Pre-Algebra?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

or Math 8

OpenStudy (anonymous):

no its pre algebra

OpenStudy (anonymous):

ohh lol Im taking algebra right now

OpenStudy (anonymous):

What grade are you in?

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

You?

OpenStudy (anonymous):

Im supposed to be in 9th but last year i started online school and throughout the whole year i was sooo lost! it was really difficult and i did pass the Staar but...they held me back because of my grades throughout the whole year :/

OpenStudy (anonymous):

Ohh thats funny cause i just started online school this year in october and i never get bad grades but it is really weird so i had all bad grades last semester

OpenStudy (anonymous):

what online school are u with?

OpenStudy (anonymous):

like all barelly passing in the 60%

OpenStudy (anonymous):

but it wasnt like i didnt know how just the school is hard to do/work

OpenStudy (anonymous):

Yeah! and its not even the actual work thats hard, its keeping up with like the progress on some of the lessons Im an A+ student in all my other sun]bjects...excpet math and literature of course! Im in k12

OpenStudy (anonymous):

*subjects

OpenStudy (anonymous):

LOL AND MEEE TO!! through what state

OpenStudy (anonymous):

Im in El paso texas!!! how bout you??

OpenStudy (anonymous):

Oh wow and im in seattle washington

OpenStudy (anonymous):

Ehh it would of been nice to know at least one person at the STAAR exams :/

OpenStudy (anonymous):

Ya..Too bad and ive never done a starr exam yet?

OpenStudy (anonymous):

i think its only a texas thing Im gonna go ask more questions so talk to you later~

OpenStudy (anonymous):

okay tag me in them!